# RSTeller: Scaling Up Visual Language Modeling in Remote Sensing with Rich Linguistic Semantics from Openly Available Data and Large Language Models

### v1

@@ 4 Experiments
    Data Scalability and Model Performance Analysis: The first experiment explores the relationship between the performance of VLMs in remote sensing and the scale of domain-specific data, specifically examining the influence of the number of RS-related image-text pairs. This experiment aims to understand how data volume affects model accuracy and generalization capabilities in open-set conditions.
@@
    Dataset Effectiveness Evaluation: We then evaluate the effectiveness of the RSTeller dataset through continual pre-training using various CLIP model checkpoints. This step assesses whether incremental learning with our dataset can enhance the model’s performance on standard benchmarks. Additionally, this experiment explores the impact of different pre-trained checkpoints on the model’s performance.
@@ 4.1 Data
The training data utilized in our experiments consists of three distinct groups: the RSTeller dataset, the SkyScript (Wang et al., 2024) dataset, and the LAION-10M dataset, which is a subset of the LAION-400M (Schuhmann et al., 2021). The detailed information is provided in Table 5 and explained below.
@@
The RSTeller dataset, which we propose, consists of approximately 2.5 million image-text pairs. There are more than two captions for one image patch in this dataset. To ensure clarity and prevent redundancy during the training of the CLIP model, each image patch appears only once in a training batch. We implement a sampling strategy where image patches for training batches are selected first, followed by a random selection of captions for these images. Both raw and augmented captions have an equal probability of 50% of being chosen for training.
The SkyScript dataset, similarly sourced from OSM and GEE, enriches our training with an additional 2.5 million image-text pairs. Each image in this dataset is uniquely associated with one caption. This dataset is integral to augmenting the RS domain knowledge and is merged with the RSTeller dataset to foster a robust study of the relation between model performance and domain data scale in our first experiment.
@@
For evaluative purposes, we employ a series of zero-shot classification and image retrieval tasks. For zero-shot classification, we use seven widely recognized and comparatively extensive benchmark datasets, including AID (Xia et al., 2017), EuroSAT (Helber et al., 2019), fMoW (Christie et al., 2018), Million-AID (Long et al., 2021), PatternNet (Zhou et al., 2018), NWPU-RESISC45 (RESISC) (Cheng et al., 2017), and RSI-CB256 (RSI-CB) (Li et al., 2020). For image retrieval evaluation, three common benchmark datasets are used, including UCM Captions (Qu et al., 2016), RSICD (Lu et al., 2017), and RSITMD (Yuan et al., 2022). Details of these benchmark datasets are presented in Tables 6 and 7, following the data usage protocols established in the prior study as SkyScript.
@@ 4.2 CLIP Continual Pre-training
Our study adopts continual pre-training to investigate the relationship between the scale of domain-specific data and performance enhancements in the CLIP model. We examine the influence of our innovative RSTeller dataset on model performance. Known challenges associated with training exclusively on new domain data include significant performance degradation due to catastrophic forgetting and potential overfitting. To mitigate these challenges, our experiments incorporate data from the original domains, specifically LAION-10M, throughout the continual pre-training process. We consistently sample one-third of the data from LAION-10M for each training batch while the rest are sampled from the RS domain. This technique is applied across all our experiments to ensure balanced training and maintain model stability.
We analyzed two versions of CLIP models: ViT-B/32 and ViT-L/14. The ViT-B configuration utilizes 12 transformer layers and 12 attention heads, whereas the ViT-L configuration comprises 24 transformer layers and 16 attention heads. The denominations /32 and /14 represent the respective patch sizes, which are $32\times 32$ and $14\times 14$ pixels, processed by the model tokenizer. Initial model setups are derived from various checkpoints, including those from the OpenCLIP project (Ilharco et al., 2021) trained on the LAION-2B dataset (Schuhmann et al., 2022), which provides a baseline of regularly pre-trained models on generic web-based image-text data. Additionally, we utilized checkpoints commonly employed by the industrial community, trained on LAION-400M (Schuhmann et al., 2021), WIT (Radford et al., 2021), and DataComp (Gadre et al., 2023), to ascertain the benefits conferred by our RSTeller dataset in the RS domain. A brief comparison of the data amount among these datasets and all the checkpoints used in our experiments are presented in Table 8. The model checkpoints are represented by the identifier used in the OpenCLIP repository (Ilharco et al., 2021).
@@
Fig. 12: Relationship between domain data scale and zero-shot classification performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on LAION-2B. Scaling domain data consistently improves the performance of the model on the validation dataset.
Fig. 13: Relationship between domain data scale and zero-shot image retrieval performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on LAION-2B. Scaling domain data consistently improves the performance of the model on the validation dataset.
@@
Experimentation was conducted using the OpenCLIP framework, incorporating automatic mixed-precision (AMP) to expedite training and optimize performance. For the experiments that scale domain data (Section 4.3), gradient checkpointing was employed to minimize memory usage, and the standard Adam optimizer was utilized with a learning rate of $1\times 10^{-5}$. Both linear warm-up and cosine annealing learning rate schedules were implemented. Each experiment involved random iterations over the data for 5,000 steps with a total batch size of 8,192 across 4 NVIDIA A800 GPUs. For the dataset effectiveness evaluation experiments (Section 4.4), gradient checkpointing was disabled, and a total batch size of 1,024 was used on 4 NVIDIA A800 GPUs for more efficient training. The training steps and learning rate were adjusted to 20,000 steps and a learning rate of $1\times 10^{-8}$, respectively, to optimize performance.
4.3 Model Performance wrt. Domain Data Scale
This experiment was designed to investigate whether scaling the domain-specific data enhances model performance. To this end, we amalgamated our RSTeller dataset with the SkyScript dataset to create a robust RS domain dataset comprising approximately 5 million image-text pairs. This extensive dataset allows for a detailed analysis of the model’s performance as it relates to variations in domain data volume. We partitioned the data into four equal subsets, each representing about one-fourth of the total data, and ensured no overlap between the subsets. In our experiments, we incrementally incorporated these subsets into the training process, thereby systematically increasing the domain data scale. Models pre-trained on LAION-2B were used as the starting point for the continual pre-training. The models were then trained on approximately 25%, 50%, 75%, and 100% of the domain data, respectively.
Most benchmarks show improvement when domain data is scaled up. Notably, the slopes of both models’ error rates are negative ($k_{Average}=-0.75$ for ViT-B/32 and $k_{Average}=-0.34$ for ViT-L/14). Benchmarks such as PatternNet, RESISC, and AID exhibit substantial improvements, as indicated by the slopes of their linear fits (e.g., $k_{PatternNet}=-1.63$, $k_{RESISC}=-1.03$ for the ViT-B/32 model, and $k_{AID}=-1.36$ for the ViT-L/14 model). These results underscore the positive impact of enlarging the domain-specific dataset on the model’s classification accuracy. The EuroSAT on the ViT-L/14 model, on the other hand, demonstrates an unexpected increase in top-1 classification error as the domain data scale rises ($k_{EuroSAT}=0.76$). This anomaly suggests potential noise in the training process or an inherent difficulty in this benchmark that requires more careful consideration in scaling the domain data.
For some benchmarks, such as AID and RSI-CB, there is an initial performance degradation when domain-specific data is first introduced. In extreme cases, even with all the domain data used for training, the fine-tuned model performance still does not surpass the baseline (e.g., AID for ViT-B/32). Nevertheless, the continual inclusion of domain data leads to consistent performance improvement. For instance, the AID benchmark shows a steady decline in error rates despite the initial degradation, as indicated by the slopes ($k_{AID}=-0.72$ for ViT-B/32 and $k_{AID}=-1.36$ for ViT-L/14). This trend suggests that further scaling of domain data could potentially surpass the baseline performance of models pre-trained on generic web-based data.
The varied slopes of the linear fits also highlight the differential impact of domain data across different benchmarks. For example, in the ViT-B/32 experiment, the significant negative slope observed for PatternNet ($k_{PatternNet}=-1.63$) compared to the minimal impact on Million-AID ($k_{Million-AID}=-0.07$) suggests that the effectiveness of domain-specific data augmentation varies across different datasets. This variation indicates inherent differences in the benchmarks themselves and possibly the domain specificity of the data they contain.
The experimental results for zero-shot image retrieval tasks, illustrated in Fig. 13, reveal trends consistent with those observed in the classification experiments. The x-axis represents the volume of RS-related domain data used for continual pre-training, while the y-axis indicates the evaluation metric, which is the average recall rates of top-1 (R@1), top-5 (R@5), and top-10 (R@10), subtracted from 100.
For both the ViT-B/32 and ViT-L/14 models, there is a general improvement in performance with increasing domain data, as indicated by the negative slopes of the linear fits ($k_{Average}=-0.33$ for ViT-B/32 and $k_{Average}=-0.57$ for ViT-L/14). This trend is evident across multiple benchmarks, highlighting the benefits of enlarging domain-specific datasets for image retrieval tasks.
However, a notable difference from the classification results is observed in the ViT-L/14 model’s performance for the retrieval task. None of the fine-tuned ViT-L/14 models were able to surpass the original checkpoint performance across all benchmarks, even with the full domain data included. Despite this, there is a consistent trend of improving recall rates with the increase in domain-specific data, suggesting that further scaling could potentially lead to superior performance in the future.
The experimental results confirm that increasing domain-specific data generally enhances model performance in the field of remote sensing. Despite initial degradation observed in specific benchmarks, continuous improvement with increased data volume underscores the potential for further advancements. These findings validate the significance of this work, highlighting the proposed automated workflow’s efficiency in generating larger datasets at a low cost. This capability offers substantial benefits to the research community, particularly for training VLMs related to remote sensing.
@@ 4.4 Dataset Effectiveness
Model Pretrained Dataset AID EuroSAT fMoW Million-AID RESISC PatternNet RSI-CB Average
ViT-B/32 LAION-2B 68.10 24.40 16.00 53.75 60.46 64.36 38.11 46.45
+RSTeller 61.15 31.11 14.70 51.36 54.36 66.37 40.85 45.70
$\Delta$6.95 6.71 -1.30 -2.39 -6.10 2.01 2.74 -0.75
LAION-400M 60.20 29.19 13.12 43.91 52.50 51.02 31.85 40.26
+RSTeller 60.25 37.93 14.25 41.14 47.36 50.37 36.90 41.17
$\Delta$ 0.05 8.74 1.12 -2.77 -5.14 -0.66 5.05 0.91
WIT 57.40 25.04 15.42 51.52 54.44 52.80 29.92 40.93
+RSTeller 59.65 36.44 15.60 52.79 55.13 62.29 35.83 45.39
$\Delta$ 2.25 11.41 0.18 1.27 0.69 9.48 5.92 4.46
DataComp 66.10 24.15 15.47 49.47 62.56 60.14 35.24 44.73
+RSTeller 63.65 36.89 15.61 49.73 58.70 64.28 41.39 47.18
$\Delta$2.45 12.74 0.14 0.26 -3.86 4.13 6.15 2.44
ViT-L/14 LAION-2B 67.80 38.04 24.39 63.02 70.42 72.79 42.92 54.20
+RSTeller 66.05 45.33 21.29 60.96 68.54 76.22 46.23 54.95
$\Delta$1.75 7.30 -3.10 -2.06 -1.88 3.42 3.31 0.75
LAION-400M 66.55 33.67 19.85 59.72 66.11 67.11 43.62 50.95
+RSTeller 65.05 47.07 17.36 56.89 62.14 69.83 41.89 51.46
$\Delta$1.50 13.41 -2.49 -2.83 -3.97 2.72 -1.73 0.52
WIT 69.05 41.70 25.48 57.72 66.62 71.32 43.05 53.56
+RSTeller 67.60 56.30 22.30 58.16 64.68 77.14 47.32 56.21
$\Delta$1.45 14.59 -3.18 0.44 -1.95 5.82 4.27 2.65
DataComp 69.65 52.56 27.57 66.89 70.98 78.22 43.59 58.49
+RSTeller 68.85 55.22 23.15 65.24 69.96 79.82 47.71 58.57
$\Delta$0.80 2.67 -4.42 -1.65 -1.02 1.60 4.13 0.07
@@
To assess the benefits of the RSTeller dataset for pre-trained models, a series of continual training experiments were conducted on various CLIP checkpoints, evaluating their zero-shot capabilities before and after the continual pre-training. Notably, in this experiment, the SkyScript data was excluded from the training data, differing from the previous experiment. The LAION-10M was also utilized to prevent catastrophic forgetting and overfitting.
@@
The evaluation results are presented in Tables 9 and 10. These tables report the top-1 accuracy for zero-shot classification tasks and the average recall rate for zero-shot text-to-image retrieval tasks. For each pre-trained checkpoint, the performance of the baseline checkpoints on the benchmark datasets, the performance after continual pre-training on the RSTeller dataset, and the corresponding improvements (marked as $\Delta$) are shown. The positive improvements are highlighted in red, while the negative changes are highlighted in blue. For each benchmark, the largest and least improvements are highlighted in boldface, and the best results within the continual pre-trained models are underlined.
@@
Table 9 summarizes the top-1 zero-shot accuracy on various benchmark datasets before and after continual pre-training on the RSTeller dataset. Except for the ViT-B/32 model pre-trained on LAION-2B, all other models show improvements in the average results (marked as $\Delta$ values in red). This suggests that the RSTeller dataset enhances the performance of the VLMs, making it a valuable resource for pre-training VLMs in the remote sensing domain.
@@
Despite the overall improvements in the average zero-shot classification results, the degree of enhancement varies across different benchmarks. EuroSAT consistently shows the most significant improvements (marked as bold $\Delta$ values in red) for all models and pre-trained checkpoints, with a maximum increase of $14.59\%$ for the ViT-L/14 model pre-trained on WIT. In contrast, some benchmark datasets such as AID, fMoW, and RESISC experienced notable decreases. However, these decreases were not consistent across all pre-trained checkpoints. For example, the ViT-B/32 model pre-trained on LAION-2B showed a significant decrease in AID by $6.95\%$, whereas the same model pre-trained on WIT saw an increase of $2.25\%$ on the same benchmark. This variability suggests that the RSTeller dataset may favor certain pre-trained checkpoints and tasks.
@@
Table 10: Zero-shot image-to-text retrieval recall rates (%) on benchmark datasets before and after continual pre-training with RSTeller dataset
Model Pretrained Dataset RSICD RSITMD UCM-Captions Average
ViT-B/32 LAION-2B 16.73 27.50 57.21 33.81
+RSTeller 17.28 27.59 54.55 33.14
$\Delta$ 0.55 0.09 -2.65 -0.67
LAION-400M 14.49 22.54 54.38 30.47
+RSTeller 15.40 24.43 53.93 31.25
$\Delta$ 0.91 1.89 -0.44 0.78
@@
+RSTeller 15.10 26.49 53.40 31.67
$\Delta$0.78 1.46 0.53 0.41
ViT-L/14 LAION-2B 20.34 31.81 62.95 38.37

### v4

@@ 4 Experiments
    Dataset Effectiveness Evaluation: We begin by evaluating the effectiveness of the RSTeller dataset through continual pre-training with various CLIP model checkpoints. This step assesses whether incremental learning with our dataset can enhance the model’s performance on standard benchmarks. Additionally, this experiment explores the impact of different pre-trained checkpoints on the model’s performance.
@@
    Ablation Study on Training Data: To assess the contribution of each component in the training data used during the continual pre-training process, we conducted an ablation study, systematically removing one component of the training data at a time. We examined the impact of excluding common knowledge, removing named entities, omitting LLM interpretation, and excluding the caption revision task. This analysis helps identify the significance of each element in enhancing model performance.
  3. 3.
    Data Scalability and Model Performance Analysis: The final experiment explores the relationship between the performance of VLMs in remote sensing and the scale of domain-specific data, specifically examining the influence of the number of RS-related image-text pairs. This experiment aims to understand how data volume affects model accuracy and generalization capabilities in open-set conditions.
@@ 4.1 Data
The training data utilized in our experiments consist of three distinct groups: the RSTeller dataset, the SkyScript (Wang et al., 2024) dataset, and the LAION-10M dataset, which is a subset of the LAION-400M (Schuhmann et al., 2021). The detailed information is provided in Table 5 and is explained below.
@@
The RSTeller dataset, which we propose, consists of approximately 2.6 million image-text pairs. There are two captions for one image patch in this dataset. To ensure clarity and prevent redundancy during the training of the CLIP model, each image patch appears only once in a training batch. We implement a sampling strategy where image patches for training batches are selected first, followed by a random selection of captions for these images.
The SkyScript dataset, similarly sourced from OSM and GEE, enriches our training with an additional 2.6 million image-text pairs. Each image in this dataset is uniquely associated with one caption. This dataset is integral to augmenting the RS domain knowledge and is merged with the RSTeller dataset to foster a robust study of the relation between model performance and domain data scale in our last experiment.
@@
For evaluative purposes, we employ a series of zero-shot classification and image retrieval tasks. For zero-shot classification, we use eight widely recognized and comparatively extensive benchmark datasets, including AID (Xia et al., 2017), EuroSAT (Helber et al., 2019), fMoW (Christie et al., 2018), Million-AID (Long et al., 2021), PatternNet (Zhou et al., 2018), NWPU-RESISC45 (RESISC) (Cheng et al., 2017), RSI-CB256 (RSI-CB) (Li et al., 2020), and ImageNet-1K (Russakovsky et al., 2015). The first seven datasets are from the RS domain, while ImageNet-1K pertains to the general domain. The inclusion of ImageNet-1K allows for an evaluation of the model’s general-purpose capabilities both before and after the continual pre-training with domain-specific remote sensing data. For image retrieval evaluation, three RS benchmark datasets are used, including UCM Captions (Qu et al., 2016), RSICD (Lu et al., 2017), and RSITMD (Yuan et al., 2022). Details of these benchmark datasets are presented in Tables 6 and 7, following the data usage protocols established in the prior study as SkyScript.
@@ 4.2 CLIP Continual Pre-training
Our study adopts continual pre-training to assess the effectiveness of the proposed dataset, evaluate its individual components, and investigate the relationship between the scale of domain-specific data and performance improvements in the CLIP model. Known challenges associated with training exclusively on new domain data include significant performance degradation due to catastrophic forgetting and potential overfitting. To mitigate these challenges, our experiments incorporate data from the original domains, specifically LAION-10M, throughout the continual pre-training process. We consistently sample one-third of the data from LAION-10M for each training batch while the rest are sampled from the RS domain. This technique is applied across all our experiments to ensure balanced training and maintain model stability. Additionally, the effectiveness of this technique is validated the second experiment.
We evaluated two versions of the CLIP model: ViT-B/32 and ViT-L/14. The ViT-B model uses 12 transformer layers and 12 attention heads, while ViT-L has 24 transformer layers and 16 attention heads. The numbers /32 and /14 refer to the patch sizes, $32\times 32$ and $14\times 14$ pixels, respectively, processed by the model tokenizer. We used the standard OpenCLIP implementation (Ilharco et al., 2021) and initialized the model with four pre-trained checkpoints: the official OpenAI CLIP pre-trained on WIT (Radford et al., 2021), models trained on LAION-400M (Schuhmann et al., 2021), DataComp (Gadre et al., 2023), and LAION-2B (Schuhmann et al., 2022). A comparison of the dataset sizes and model checkpoints used is provided in Table 8. The model checkpoints are represented by the identifier used in the OpenCLIP repository.
@@
Experiments were conducted using the OpenCLIP framework, with automatic mixed-precision (AMP) employed to accelerate training and optimize performance. The standard Adam optimizer was used, and both linear warm-up and cosine annealing learning rate schedules were implemented. In the first two experiments, which explore the effectiveness of the RSTeller dataset and perform ablation studies on the components of the training data (Section 4.3 and Section 4.4), a moderate training configuration was applied to balance the preservation of the pre-trained model feature space with the incorporation of the training data. A learning rate of $1\times 10^{-8}$ was adopted, and each experiment consisted of 20,000 steps with random data iterations, using a total batch size of 1,024 distributed across 4 NVIDIA A800 GPUs. For the final experiment (Section 4.5), which investigates the scaling of domain data on model performance, a more aggressive training configuration was used, with a learning rate of $1\times 10^{-5}$ and a total batch size of 8,192 across 4 NVIDIA A800 GPUs, running for 5,000 steps. Gradient checkpointing was employed to minimize memory usage. In the moderate configuration, the model does not fully overfit to the training data, while in the aggressive configuration, the model is forced to fit more closely to the training data.
@@ 4.3 Dataset Effectiveness
  Model Pre-trained Dataset AID EuroSAT fMoW Million-AID RESISC PatternNet RSI-CB Average ImageNet-1K
  ViT-B/32 LAION 2B 68.10 24.40 16.00 53.75 60.46 64.36 38.11 46.45 65.84
+RSTeller 61.50 29.56 14.79 50.98 55.16 65.71 42.78 45.78 65.30
$\Delta$6.60 5.16 -1.21 -2.77 -5.30 1.35 4.67 -0.67 -0.53
LAION 400M 60.20 29.19 13.12 43.91 52.50 51.02 31.85 40.26 59.64
+RSTeller 60.50 36.96 14.92 43.84 47.31 53.63 40.43 42.51 60.08
$\Delta$ 0.30 7.78 1.80 -0.07 -5.18 2.60 8.58 2.26 0.44
WIT 57.40 25.04 15.42 51.52 54.44 52.80 29.92 40.93 63.37
+RSTeller 60.60 36.41 15.22 52.06 56.17 62.11 38.21 45.82 63.59
$\Delta$ 3.20 11.37 -0.19 0.54 1.72 9.31 8.29 4.89 0.22
DataComp 66.10 24.15 15.47 49.47 62.56 60.14 35.24 44.73 68.98
+RSTeller 63.60 38.37 16.06 48.11 58.47 64.72 41.31 47.23 68.41
$\Delta$2.50 14.22 0.59 -1.36 -4.10 4.58 6.07 2.50 -0.57
  ViT-L/14 LAION 2B 67.80 38.04 24.39 63.02 70.42 72.79 42.92 54.20 75.24
+RSTeller 66.05 44.70 22.86 61.00 69.36 75.08 46.64 55.10 74.93
$\Delta$1.75 6.67 -1.52 -2.02 -1.06 2.29 3.72 0.90 -0.31
LAION 400M 66.55 33.67 19.85 59.72 66.11 67.11 43.62 50.95 72.77
+RSTeller 64.95 47.70 17.29 57.82 61.89 70.11 43.99 51.96 71.97
$\Delta$1.60 14.04 -2.56 -1.90 -4.22 3.00 0.37 1.02 -0.81
WIT 69.05 41.70 25.48 57.72 66.62 71.32 43.05 53.56 75.54
+RSTeller 68.95 58.78 23.31 58.51 66.61 78.31 48.96 57.63 75.69
$\Delta$0.10 17.07 -2.17 0.79 -0.01 6.99 5.91 4.07 0.16
DataComp 69.65 52.56 27.57 66.89 70.98 78.22 43.59 58.49 79.19
+RSTeller 66.75 56.41 24.42 63.97 69.62 78.85 45.75 57.97 79.10
$\Delta$2.90 3.85 -3.15 -2.92 -1.36 0.63 2.17 -0.53 -0.09
@@
To assess the benefits of the RSTeller dataset for pre-trained models, a series of continual training experiments were conducted on various CLIP checkpoints, evaluating their zero-shot capabilities before and after the continual pre-training. Notably, in this experiment, LAION-10M was utilized to prevent catastrophic forgetting and overfitting.
@@
The evaluation results are presented in Tables 9 and 10. These tables report the top-1 accuracy for zero-shot classification tasks, as well as the average recall rates at top-1 (R@1), top-5 (R@5), and top-10 (R@10) for zero-shot text-to-image retrieval tasks. For each pre-trained checkpoint, the performance of the baseline checkpoints on the benchmark datasets, the performance after continual pre-training on the RSTeller dataset, and the corresponding improvements (marked as $\Delta$) are shown. To provide a comprehensive overview of the model’s adaptability, the results from seven remote sensing datasets are averaged to reflect overall performance changes in the remote sensing domain, while results on ImageNet-1K are separately reported to assess the model’s robustness against catastrophic forgetting and overfitting. The positive improvements are highlighted in blue, while the negative changes are highlighted in red. For each benchmark, the largest and least improvements are highlighted in boldface, and the best results within the continual pre-trained models are underlined.
@@
Table 9 presents the top-1 zero-shot accuracy on various benchmark datasets before and after continual pre-training on the RSTeller dataset. Except for the ViT-B/32 model pre-trained on LAION-2B and the ViT-L/14 model pre-trained on DataComp, which exhibit a minor decline in performance on the remote sensing domain, all other models show improvements in the average results (marked as $\Delta$ values in blue). This suggests that the RSTeller dataset enhances the performance of the VLMs, making it a valuable resource for pre-training VLMs in the remote sensing domain. Additionally, all models exhibit minimal performance changes when evaluated on the ImageNet-1K dataset, with most experiencing a slight decline. However, some models, such as ViT-L/14 and ViT-B/32 trained on WIT, show performance gains on ImageNet-1K. This improvement may be attributed to the inclusion of common-domain data during the training process, which positively impacts generalization.
@@
Despite the overall improvements in the average zero-shot classification results, the degree of enhancement varies across different benchmarks. EuroSAT consistently shows the most significant improvements for almost all models and pre-trained checkpoints, with a maximum increase of $17.07\%$ for the ViT-L/14 model pre-trained on WIT. In contrast, some benchmark datasets such as AID, fMoW, and RESISC experienced notable decreases. However, these decreases were not consistent across all pre-trained checkpoints. For example, the ViT-B/32 model pre-trained on LAION-2B showed a significant decrease in AID by $6.60\%$, whereas the same model pre-trained on WIT saw an increase of $3.20\%$ on the same benchmark. This variability suggests that the RSTeller dataset may favor certain pre-trained checkpoints and tasks.
@@
Table 10: Zero-shot text-to-image retrieval recall rates (%) on benchmark datasets before and after continual pre-training with RSTeller dataset
Model Pre-trained Dataset RSICD RSITMD UCM-Captions Average
ViT-B/32 LAION 2B 16.73 27.50 57.21 33.81
+RSTeller 16.90 27.04 52.52 32.15
$\Delta$ 0.17 -0.46 -4.69 -1.66
LAION 400M 14.49 22.54 54.38 30.47
+RSTeller 14.07 22.54 53.49 30.03
$\Delta$0.43 0.00 -0.88 -0.44
@@
+RSTeller 14.11 24.59 51.90 30.20
$\Delta$1.77 -0.44 -0.97 -1.06
ViT-L/14 LAION 2B 20.34 31.81 62.95 38.37
@@
Table 11: Ablation study on zero-shot image classification and text-to-image retrieval tasks.
     Settings Zero-shot Image Classification Zero-shot text-to-image retrieval
AID EuroSAT fMoW Million-AID RESISC PatternNet RSI-CB Average ImageNet-1K RSICD RSITMD UCM-Captions Average
  Baseline 68.95 58.78 23.31 58.51 66.61 78.31 48.96 57.63 75.69 16.56 28.66 59.50 34.91
w/o LAION10M -1.30 0.89 -0.44 -0.24 -1.89 -1.31 -2.38 -0.95 -0.90 -0.05 -0.36 -0.00 -0.14
w/o named entities -0.00 0.04 -0.03 -0.05 -0.03 -0.08 -0.08 -0.03 -0.00 -0.07 0.09 -0.00 0.01
w/o LLM interpretation -1.85 -3.37 1.29 -0.94 -0.57 -1.05 -2.32 -1.26 -0.01 0.42 0.05 -1.95 -0.49
w/o Task 3 -0.15 -0.37 -0.20 -0.57 -0.73 -0.20 -0.48 -0.39 -0.05 -0.64 -0.28 0.27 -0.22